OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!! Please help me to simplify this rational expression - ax+bx a+b -------- / ----- x^2 2x **Note the / means to divide

satellite73 (satellite73):

\[\frac{ax+bx}{x^2}\times \frac{2x}{a+b}\] is step one

satellite73 (satellite73):

then \[\frac{x(a+b)\times 2x}{x^2(a+b)}\] then some cancellation

OpenStudy (abbycross167):

So the a+b would cancel out then what?

OpenStudy (abbycross167):

x*2x ----- x^2

OpenStudy (mathmale):

Can your latest result be cancelled further? Try it. Simplify x*2x.

jhonyy9 (jhonyy9):

so bc. x*2x=2x^2 in this way x^2 will cancel out from the top and bottom too and will remain there jus the 2